The equation of two progressive waves is given by y1 = Asin(kx-wt) and y2 = Asin(kx -wt+θ). Show that
a) y= 2A sin(kx-wt)
b) y= 0
1
Expert's answer
2020-01-20T05:21:29-0500

As per the question,


equation of progressive waves are "y_1 = A \sin(kx-wt)" and "y_2 = A \sin(kx -wt+θ)."

When two waves are getting interfere to each other, then

"Y=y_1+y_2"

"Y=A \sin(kx-wt)+A \sin(kx -wt+θ)."

"Y=A(2\sin(\dfrac{kx-wt+kx-wt+\theta}{2})\cos(\dfrac{kx-wt-kx+wt+\theta}{2}))"

"Y=A(2\sin(\dfrac{2(kx-wt)+\theta)}{2})\cos(\dfrac{\theta}{2}))"

"Y=A(2\sin((kx-wt)+\dfrac{\theta}{2})\cos(\dfrac{\theta}{2}))"

a) If "\theta=0^\circ"

Then "Y=2A\sin(kx-wt)"

b) If "\theta=\pi"

Y=0